Theatre contains 500 seats and the ticket prices for the recent plays were $43 or adults and $28 or children. Total proceeds were $16,805. How many of each were sold?
Answer by Cromlix(4381)

Adults = x
Children = y
x + y = 500.........1
43x + 28y = 16805......2
Multiply (1) by 28
28x + 28y = 14000.....3
43x + 28y = 16805.....2
Subtract (3) from (2)
15x = 2805
x = 187
Substitute x = 187 into Eq (1)
x + y = 500
187 + y = 500
y = 500 - 187
y = 313
Therefore 187 Adult's tickets sold
313 Children's tickets sold
Hope this helps.
:-)